# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ VIII-A2 Optical Wireless Communications
    Given the aforementioned benefits, OWC is applicable to a wide range of scenarios, such as short-range cellular access, cellular backhaul links, hybrid radio-optical networks, D2D/M2M/V2V communications, optical IoT, positioning and sensing, intra-device connections, on-board communications, short-range indoor communications, long-range outdoor point-to-point communications, space communications, underwater communications, and underground communications [242, 258, 252]. Different OWC technologies have their own characteristics and the corresponding advantages which differ from one to another. While VLC supports fast, low-cost, and energy-efficient short-range communication in indoor environments, FSO provides ultra-high-speed, highly directive, and low-cost long-range point-to-point outdoor connectivity. Consequently, VLC and FSO are promising technologies to complement 6G networks, with the main application scenarios being short-range indoor access and long-range outdoor backhaul links, respectively.
    There are many challenges to overcome before freeing the potential of OWC and integrating it into future mobile networks as a complementary technology. Each OWC technology has its own set of unique challenges. For VLC, the main fundamental challenges include limited communication range, easy blockage, ambient light interference, line-of-sight (LOS) requirement, random receiver orientation, uplink communication, mobility, and the non-linearity of devices [254, 243, 258]. Hybrid RF-VLC networks have been recognized as a promising solution to tackle many of these challenges [259, 260, 261]. In general, hybrid systems consist of two or more networks that can be accessed in different ways based on the network and application types [259]. There are many possible access types, such as accessing the strongest network, accessing both networks simultaneously, accessing one network for downlink and the other for uplink, access based on the quality of service (QoS) requirements, and access the highest priority network and the other is backup [259]. The design of hybrid networks is highly complicated, with many challenging tasks, such as network selection, network access, handover, load balancing, and resource allocation [259]. AI/ML is a promising tool to assist in many of these complex tasks. For example, AI/ML can be applied to resource management and traffic prediction problems. More details on the hybrid RF-VLC systems are available in [259, 260, 261].
@@
    Pointing errors can be mitigated by pointing the transmitter toward the receiver to obtain an LOS condition, acquiring the transmitted signal at the receiving end by aligning the receiver in the direction of the beam, and maintaining accurate pointing and acquisition via adaptive tracking with appropriate measurement and feedback mechanisms. Further details on pointing, acquisition, and tracking techniques can be found in [252]. Diversity methods aim to improve the link reliability by transmitting multiple copies of the same information. Three main diversity dimensions are time, frequency, and space. At a high level, the objective of hybrid RF-FSO systems is to eliminate common disadvantages while exploiting the benefits of both technologies. As discussed earlier, hybrid networks introduce many new challenges that must be tackled before successful deployment. Further details regarding the challenges of hybrid RF-FSO systems are presented in [262].
@@
  • Literature and Future Directions: Recent survey papers have reviewed the fundamentals, latest advances, open problems, and main literature of OWC [242, 263, 264, 259, 265, 266, 267, 268, 269, 270, 271], VLC [243, 258, 272, 273, 274, 260, 275, 276, 261, 277, 278], and FSO [244, 249, 279, 251, 280, 252, 281, 262, 282]. The study in [242] presented a comparative overview of OWC technologies by discussing classification, spectrum use, system architectures, potential applications, and open issues. In [263], a comprehensive survey was conducted on OWC channel measurements and models, focusing on indoor, outdoor, underground, and underwater communication environments. The role of OWC for 5G/6G and IoT was discussed in [264]. In [259], the authors surveyed hybrid RF-optical networks. The paper [265] presented a survey on OWC technologies, with the main focus on classification, enabling technologies, link design, mitigation of impairments, security issues, and future challenges. The authors of [266] discussed a roadmap for OWC research toward 6G deployments. The work [267] provided a thorough survey of OWC for four IoT domains: terrestrial, underwater, biomedical, and underground. In [268], the security of OWC was considered. The survey [269] reviewed the latest advancements in OWC in the context of 6G and WiFi. In [270], the paper focused on optical wireless and THz communications for 6G. The authors of [271] discussed the present and future of OWC and FSO toward the 6G era.
@@
    In [243], the authors provided a comprehensive review of VLC, covering the fundamentals, communication architectures, PHY/MAC layer designs, research platforms, applications, challenges, and future directions. In [258], VLC was studied in the context of 6G. VLC PHY layer security was discussed from the information-theoretic and signal processing perspectives in [272]. Vehicular VLC was explored in [273], reviewing state-of-the-art and identifying open problems. The article [274] surveyed the VLC system technology in terms of devices, architectures, and applications. In [260], hybrid RF-VLC systems were discussed from the perspectives of network topology, performance evaluation, potential applications, and future challenges. VLC was reviewed for IoT in [275]. The work [276] examined the design of federated learning-enabled VLC for 6G networks. In [261], the authors focused on the access and applications of hybrid VLC-RF systems. A thorough review of MIMO-VLC was provided in [277]. The tutorial in [278] discussed visible light indoor positioning.
@@
    The paper [244] introduced a classification framework for FSO links and systems. In [249], the authors discussed a research roadmap toward next-generation FSO from the perspective of system design. The work [279] studied FSO communication for indoor and outdoor application scenarios. In [251], FSO was reviewed in terms of single- and multi-beam systems, the effect of certain weather conditions, and the scalability of an FSO network. The paper [280] focused on intelligent reflecting surface (IRS)-aided FSO communication systems. In [252], a comprehensive survey was given on FSO communication, with the main emphasis on application scenarios, link reliability, mitigation methods, and multi-user systems. The authors in [281] examined a field-programmable gate array-based prototype for 6G long-range FSO communication. The study [262] reviewed hybrid FSO-RF systems in terms of switching methods, routing protocols, modulation schemes, research projects, applications, challenges, and potential solutions. In [282], a thorough survey of FSO systems was provided.
@@
    TABLE V: Summary of antenna system technologies for 6G
@@
VIII-B Antenna System Technologies for 6G
In the 6G era, antenna systems are expected to be pushed to their limits in the forms of ultra-massive MIMO, RISs, and holographic MIMO (HMIMO). Ultra-massive MIMO is a promising technology to enable THz communication in cellular environments by combating severe propagation losses through high beamforming gains. An RIS is a revolutionary technology to control a wireless environment in a desirable way. HMIMO has the potential to provide very high spatial multiplexing gains due to efficient utilization of spatial dimension. These three technologies are reviewed below and summarized in Table V.
@@ VIII-B Antenna System Technologies for 6G
VIII-B1 Ultra-Massive MIMO
@@
  • Past and Present: The theoretical concept of massive MIMO was developed in 2010 for a multi-user multiple-input single-output (MISO) system with the number of antennas approaching infinity [286]. Since then, massive MIMO has been widely studied in the literature [287]. The first practical implementations of massive MIMO were witnessed in the early phase of 5G in the late 2010s, relying on antenna arrays with 64 dual-polarized elements and fully digital transceiver chains [287]. Since then, the number of antennas in practical implementations has been increasing. In the literature, the first studies on ultra-massive MIMO were conducted in the mid-2010s [285]. Currently, ultra-massive MIMO is considered as an extension of 5G mmWave massive MIMO to 6G THz frequencies, with potential applications in diverse scenarios, ranging from ultra-broadband cellular access to wireless backhauling, sensing, and localization [210].
@@
  • Literature and Future Directions: While massive MIMO theory has been well studied in the literature, the main research focus of ultra-massive MIMO is on practical designs/implementations at THz frequencies to overcome fundamental challenges, provide extreme performance, and expand the capabilities of 6G networks. Recent survey papers have reviewed the latest literature, advancements, and open problems [210, 290, 227, 291, 292, 293, 294, 289, 295]. In [210], THz ultra-massive MIMO was reviewed in terms of transceiver design, channel modeling, research advancements, and future directions. In [290], the authors studied active dynamic metasurface antennas for transmission and reception in 6G extreme massive MIMO systems. The focus was on hardware architectures, transceiver design, implementation challenges, and open research problems. Recommended future directions for metasurface-based beamforming include frequency-selective beamforming, channel estimation, hybrid passive/active metasurfaces, use cases, and experiments.
    The article [227] focused on beamforming technologies for THz ultra-massive MIMO. The covered topics include transceiver architectures, beamforming design principles, wideband beamforming, RIS-assisted beamforming, existing array technologies, and emerging applications. In [291], the authors reviewed recent massive MIMO trends toward 6G in terms of localization and sensing, AI, and non-terrestrial communications. The work [292] explored near-field communications for 6G using extremely large-scale antenna arrays. In [293], extreme MIMO was reviewed, with a focus on the hardware design. The article [294] provided a comprehensive survey on XL-MIMO for 6G, discussing hardware architectures, signal processing, and channel modeling. In [289], the authors explored 6G ultra-massive MIMO in terms of spatial multiplexing, degrees of freedom, electromagnetic characteristics, and signal processing. The authors of [295] discussed 6G XL-MIMO in the context of near-field communications.
@@
VIII-B2 Reconfigurable Intelligent Surfaces
@@
    In addition to typical passive RISs, active and hybrid RIS concepts also exist. The concept of active RIS refers to the RIS’s capability to amplify reflected signals to overcome the so-called double fading effect of a typical RIS environment [298]. Compared with passive RISs, active RISs can offer improved performance at the cost of increased power consumption [298]. While active RISs do not necessarily require active RF components, there exist hybrid architectures that rely on the combinations of passive RIS elements and some additional active RF chains [298]. Among these three different RIS concepts, there is a performance-energy/complexity/cost trade-off to be tackled to find preferable implementations for different application scenarios.
@@
  • Past and Present: Since the early 2010s, the concept of RIS has been studied under many different names, such as reconfigurable reflectarrays, programmable metasurfaces, software-controlled metasurfaces, and IRS. The term RIS was introduced in the late 2010s. At present, RIS and IRS are the most popular names, with RIS being dominant. The RIS concept was extended from passive to active (without RF chains) and hybrid versions in the early 2020s [298]. Generally, many aspects of RIS technology have been explored, ranging from fundamentals and algorithmic designs to prototyping and experiments [302, 303]. Recently, a special focus of research has been directed toward RIS-assisted 6G networks with a wide range of application scenarios [297, 218, 304, 305, 306].
@@
    Although RIS technology has received significant worldwide attention in academia and industry, standards exist only at regional levels [303]. However, coordinated efforts toward common understanding and guidelines have been taken in terms of large-scale projects and new interest and specification groups organized by international standardization bodies. For example, the EU’s Horizon 2020 program has recently funded many RIS-dedicated projects [303]. Moreover, the IEEE Communications Society has established two interest groups and an emerging technology initiative focusing on RISs to promote multidisciplinary research and international collaborations, including academic and industrial players [303]. In 2021, ETSI founded an industry specification group on RISs, with a wide range of partners from vendors and operators to research institutes and universities [303]. The expected outcome consists of white papers, technical reports, and proof-of-concepts. Further details on the industrial and standardization aspects of RISs are presented in [303].
@@ VIII-B3 Holographic MIMO
    TABLE VI: Summary of transmission scheme technologies for 6G
@@
  • Literature and Future Directions: Since the late 2010s, many aspects of active HMIMOS have been researched. The main topics include fundamental theory, channel modeling, algorithm design, practical implementation, experiments, and sensing/localization. Since HMIMOS is a new research area, all the aforementioned directions need major research and development efforts in the near future to attain an adequate level of maturity for practical implementation.
@@
VIII-C Transmission Scheme Technologies for 6G
The role of a well-designed transmission scheme is crucial for 6G since it determines how efficiently the available spectral resources can be used. In other words, the 6G transmission scheme plays a key role in turning a massive spectrum range into extreme performance to support stringent requirements and a wide range of demanding usage scenarios. To this end, 6G requires an ultra-flexible transmission scheme with novel technological elements. Some potential technologies include highly tunable multi-waveform scheme, advanced modulation and coding methods, efficient NOMA, and fast grant-free medium access, as summarized in Table VI. In the following, we provide a detailed discussion on each of these technologies.
VIII-C1 Multi-Waveform Scheme
@@
  • Opportunities and Challenges: The potential of a multi-waveform design, with highly tunable numerology, is to provide an efficient core to an ultra-flexible transmission structure to support a vast range of possible 6G frequencies, requirements, use cases, and application scenarios. Due to 6G’s massive spectrum range, relying solely on a single waveform may not be a feasible solution. A preferable approach is to select a set of waveforms to support all types of 6G scenarios. To this end, the main challenge is to develop a proper set of waveforms to support a massive spectrum range from sub-6 GHz to THz frequencies, and also to find a comprehensive set of parameters for those waveforms and the associated frame structure.
@@
  • Literature and Future Directions: In the literature, a vast variety of single-carrier, multi-carrier, and OFDM-based waveforms have been studied for 5G and beyond [338, 339, 337, 340]. Typical single-carrier waveform candidates include different variations of DFT-s-OFDM [341, 216, 337]. In the multi-carrier domain, popular waveforms include FBMC, UFMC, GFDM, and orthogonal time frequency space modulation [342, 343, 339, 344]. Numerous OFDM variants have been developed, such as F-OFDM, block-filtered OFDM, W-OFDM, cyclic postfix windowing-OFDM, time-interleaved block-windowed burst OFDM, and weighted overlap and add OFDM [338]. Since CP-OFDM has established its status in mobile networks, competing technologies aim to overcome its shortcomings. However, considering individual waveforms, each of these technologies has its own pros and cons, but there is no single method that fits well to all 6G scenarios. Consequently, multiple waveforms must be utilized in ultra-heterogeneous 6G networks.
@@
VIII-C2 Advanced Modulation and Coding Methods

### v3

@@ IX-A2 OPTICAL WIRELESS COMMUNICATIONS
    Given the aforementioned benefits, OWC is applicable to a wide range of scenarios, such as short-range cellular access, cellular backhaul links, hybrid radio-optical networks, D2D/M2M/V2V communications, optical IoT, positioning and sensing, intra-device connections, on-board communications, short-range indoor communications, long-range outdoor point-to-point communications, space communications, underwater communications, and underground communications [259, 275, 269]. Different OWC technologies have their own characteristics and the corresponding advantages which differ from one to another. While VLC supports fast, low-cost, and energy-efficient short-range communication in indoor environments, FSO provides ultra-high-speed, highly directive, and low-cost long-range point-to-point outdoor connectivity. Consequently, VLC and FSO are promising technologies to complement 6G networks, with the main application scenarios being short-range indoor access and long-range outdoor backhaul links, as illustrated in Figure 14.
    Figure 14: Main applications of visible light communication and free-space optical.
    There are many challenges to overcome before freeing the potential of OWC and integrating it into future mobile networks as a complementary technology. Each OWC technology has its own set of unique challenges. For VLC, the main fundamental challenges include limited communication range, easy blockage, ambient light interference, line-of-sight (LOS) requirement, random receiver orientation, uplink communication, mobility, and the non-linearity of devices [271, 260, 275]. Hybrid RF-VLC networks have been recognized as a promising solution to tackle many of these challenges [276, 277, 278]. In general, hybrid systems consist of two or more networks that can be accessed in different ways based on the network and application types [276]. There are many possible access types, such as accessing the strongest network, accessing both networks simultaneously, accessing one network for downlink and the other for uplink, access based on the quality of service (QoS) requirements, and access the highest priority network and the other is backup [276]. The design of hybrid networks is highly complicated, with many challenging tasks, such as network selection, network access, handover, load balancing, and resource allocation [276]. AI/ML is a promising tool to assist in many of these complex tasks. For example, AI/ML can be applied to resource management and traffic prediction problems. More details on the hybrid RF-VLC systems can be found in [276, 277, 278].
@@
    Pointing errors can be mitigated by pointing the transmitter toward the receiver to obtain an LOS condition, acquiring the transmitted signal at the receiving end by aligning the receiver in the direction of the beam, and maintaining accurate pointing and acquisition via adaptive tracking with appropriate measurement and feedback mechanisms. Further details on pointing, acquisition, and tracking techniques can be found in [269]. Diversity methods aim to improve the link reliability by transmitting multiple copies of the same information. Three main diversity dimensions are time, frequency, and space. At a high level, the objective of hybrid RF-FSO systems is to eliminate common disadvantages while exploiting the benefits of both technologies. As discussed earlier, hybrid networks introduce many new challenges that must be tackled before successful deployment. Further details regarding the challenges of hybrid RF-FSO systems can be found in [279].
@@
  • Literature and Future Directions: Recent survey papers have reviewed the fundamentals, latest advances, open problems, and main literature of OWC [276, 280, 281, 282, 283, 284, 285, 286], VLC [275, 287, 288, 289, 277, 290, 291, 278, 292, 293], and FSO [266, 294, 268, 295, 269, 296, 279, 297]. In [276], the authors surveyed hybrid RF-optical networks. The paper [280] presented a survey on OWC technologies, with the main focus on classification, enabling technologies, link design, mitigation of impairments, security issues, and future challenges. The authors of [281] discussed a roadmap for OWC research toward 6G deployments. The work [282] provided a thorough survey of OWC for four IoT domains: terrestrial, underwater, biomedical, and underground. In [283], the security of OWC was considered. The survey [284] reviewed the latest advancements in OWC in the context of 6G and WiFi. In [285], the paper focused on optical wireless and THz communications for 6G. The authors of [286] discussed the present and future of OWC and FSO toward the 6G era.
@@
    In [275], VLC was studied in the context of 6G. VLC PHY layer security was discussed from the information-theoretic and signal processing perspectives in [287]. Vehicular VLC was explored in [288], reviewing state-of-the-art and identifying open problems. The article [289] surveyed the VLC system technology in terms of devices, architectures, and applications. In [277], hybrid RF-VLC systems were discussed from the perspectives of network topology, performance evaluation, potential applications, and future challenges. VLC was reviewed for IoT in [290]. The work [291] examined the design of federated learning-enabled VLC for 6G networks. In [278], the authors focused on the access and applications of hybrid VLC-RF systems. A thorough review of MIMO-VLC was provided in [292]. The tutorial in [293] discussed visible light indoor positioning.
@@
    In [266], the authors discussed a research roadmap toward next-generation FSO from the perspective of system design. The work [294] studied FSO communication for indoor and outdoor application scenarios. In [268], FSO was reviewed in terms of single- and multi-beam systems, the effect of certain weather conditions, and the scalability of an FSO network. The paper [295] focused on intelligent reflecting surface (IRS)-aided FSO communication systems. In [269], a comprehensive survey was given on FSO communication, with the main emphasis on application scenarios, link reliability, mitigation methods, and multi-user systems. The authors in [296] examined a field-programmable gate array-based prototype for 6G long-range FSO communication. The study [279] reviewed hybrid FSO-RF systems in terms of switching methods, routing protocols, modulation schemes, research projects, applications, challenges, and potential solutions. In [297], a thorough survey of FSO systems was provided.
@@
    TABLE VII: Summary of antenna system technologies for 6G
@@
IX-B ANTENNA SYSTEM TECHNOLOGIES FOR 6G
In the 6G era, antenna systems are expected to be pushed to their limits in the forms of ultra-massive MIMO, RISs, and holographic MIMO (HMIMO). Ultra-massive MIMO is a promising technology to enable THz communication in cellular environments by combating severe propagation losses through high beamforming gains. An RIS is a revolutionary technology to control a wireless environment in a desirable way. HMIMO has the potential to provide very high spatial multiplexing gains due to efficient utilization of spatial dimension. These three technologies are reviewed below and summarized in Table VII.
@@ IX-B ANTENNA SYSTEM TECHNOLOGIES FOR 6G
IX-B1 ULTRA-MASSIVE MIMO
@@
  • Past and Present: The theoretical concept of massive MIMO was developed in 2010 for a multi-user multiple-input single-output (MISO) system with the number of antennas approaching infinity [302]. Since then, massive MIMO has been widely studied in the literature [303]. The first practical implementations of massive MIMO were witnessed in the early phase of 5G in the late 2010s, relying on antenna arrays with 64 dual-polarized elements and fully digital transceiver chains [303]. Since then, the number of antennas in practical implementations has been steadily increasing, reaching already several hundreds [99]. In the literature, the first studies on ultra-massive MIMO were conducted in the mid-2010s [301]. Currently, ultra-massive MIMO is considered as an extension of 5G mmWave massive MIMO to 6G THz frequencies, with potential applications in diverse scenarios, ranging from ultra-broadband cellular access to wireless backhauling, sensing, and localization [229].
@@
  • Literature and Future Directions: While massive MIMO theory has been well studied in the literature, the main research focus of ultra-massive MIMO is on practical designs/implementations at THz frequencies to overcome fundamental challenges, provide extreme performance, and expand the capabilities of 6G networks. Recent survey papers have reviewed the latest literature, advancements, and open problems [229, 306, 244, 307, 308, 309, 310, 305, 311]. In [229], THz ultra-massive MIMO was reviewed in terms of transceiver design, channel modeling, research advancements, and future directions. In [306], the authors studied active dynamic metasurface antennas for transmission and reception in 6G extreme massive MIMO systems. The focus was on hardware architectures, transceiver design, implementation challenges, and open research problems. Recommended future directions for metasurface-based beamforming included frequency-selective beamforming, channel estimation, hybrid passive/active metasurfaces, use cases, and experiments.
    The article [244] focused on beamforming technologies for THz ultra-massive MIMO. The covered topics included transceiver architectures, beamforming design principles, wideband beamforming, RIS-assisted beamforming, existing array technologies, and emerging applications. In [307], the authors reviewed recent massive MIMO trends toward 6G in terms of localization and sensing, AI, and non-terrestrial communications. The work [308] explored near-field communications for 6G using extremely large-scale antenna arrays. In [309], extreme MIMO was reviewed, with a focus on the hardware design. The article [310] provided a comprehensive survey on XL-MIMO for 6G, discussing hardware architectures, signal processing, and channel modeling. In [305], the authors explored 6G ultra-massive MIMO in terms of spatial multiplexing, degrees of freedom, electromagnetic characteristics, and signal processing. The authors of [311] discussed 6G XL-MIMO in the context of near-field communications.
@@
IX-B2 RECONFIGURABLE INTELLIGENT SURFACES
@@
    In addition to typical passive RISs, active and hybrid RIS concepts also exist. The concept of active RIS refers to the capability to amplify reflected signals to overcome the so-called double fading effect of a typical RIS environment [314]. Compared with passive RISs, active RISs can offer improved performance at the cost of increased power consumption [314]. While active RISs do not necessarily require active RF components, there exist hybrid architectures that rely on the combinations of passive RIS elements and some additional active RF chains [314]. Among these three different RIS concepts, there is a performance-energy/complexity/cost trade-off to be tackled to find preferable implementations for different application scenarios.
@@
  • Past and Present: Since the early 2010s, the concept of RIS has been studied under many different names, such as reconfigurable reflectarrays, programmable metasurfaces, software-controlled metasurfaces, and IRS. The term RIS was introduced in the late 2010s. At present, RIS and IRS are the most popular names, with RIS being dominant. The RIS concept was extended from passive to active (without RF chains) and hybrid versions in the early 2020s [314]. Generally, many aspects of RIS technology have been explored, ranging from fundamentals and algorithmic designs to prototyping and experiments [318, 319]. Recently, a special focus of research has been directed toward RIS-assisted 6G networks, with a wide range of application scenarios [313, 237, 320, 321, 322].
@@
    Although RIS technology has received significant worldwide attention in academia and industry, standards exist only at regional levels [319]. However, coordinated efforts toward common understanding and guidelines have been taken in terms of large-scale projects and new interest and specification groups organized by international standardization bodies. For example, the European Horizon 2020 program has recently funded many RIS-dedicated projects [319]. Moreover, the IEEE Communications Society has established two interest groups and an emerging technology initiative focusing on RISs to promote multidisciplinary research and international collaborations, including academic and industrial players [319]. In 2021, ETSI founded an industry specification group on RISs, with a wide range of partners from vendors and operators to research institutes and universities [319]. The expected outcome consists of white papers, technical reports, and proof-of-concepts. Further details on the industrial and standardization aspects of RISs can be found in [319].
@@ IX-B3 HOLOGRAPHIC MIMO
    TABLE VIII: Summary of transmission scheme technologies for 6G
@@
  • Literature and Future Directions: Since the late 2010s, many aspects of active HMIMOS have been researched. The main topics included fundamental theory, channel modeling, algorithm design, practical implementation, experiments, and sensing/localization. Since HMIMOS is a new research area, all the aforementioned directions need major research and development efforts in the near future to attain an adequate level of maturity for practical implementation.
@@
IX-C TRANSMISSION SCHEME TECHNOLOGIES FOR 6G
The role of a well-designed transmission scheme is crucial for 6G since it determines how efficiently the available spectral resources can be used. In other words, the 6G transmission scheme plays a key role in turning a massive spectrum range into extreme performance to support stringent requirements and a wide range of demanding usage scenarios. To this end, 6G requires an ultra-flexible transmission scheme with novel technological elements. Some potential technologies include highly tunable multi-waveform scheme, advanced modulation and coding methods, efficient NOMA, and fast grant-free medium access, as summarized in Table VIII. In the following, we provide a detailed discussion on each of these technologies.
IX-C1 MULTI-WAVEFORM SCHEME
@@
  • Opportunities and Challenges: The potential of a multi-waveform design, with highly tunable numerology, is to provide an efficient core to an ultra-flexible transmission structure to support a vast range of possible 6G frequencies, requirements, use cases, and application scenarios. Due to the massive spectrum range of 6G, relying solely on a single waveform may not be a feasible solution. A preferable approach is to select a set of waveforms to support all types of 6G scenarios. To this end, the main challenge is to develop a proper set of waveforms to support a massive spectrum range from sub-6 GHz to THz frequencies, and also to find a comprehensive set of parameters for those waveforms and the associated frame structure.
@@
  • Literature and Future Directions: In the literature, a vast variety of single-carrier, multi-carrier, and OFDM-based waveforms has been studied for 5G and beyond [354, 355, 353, 356]. Typical single-carrier waveform candidates include different variations of DFT-s-OFDM [357, 235, 353]. In the multi-carrier domain, popular waveforms include FBMC, UFMC, GFDM, and orthogonal time frequency space modulation [358, 359, 355, 360]. Numerous OFDM variants have been developed, such as F-OFDM, block-filtered OFDM, W-OFDM, cyclic postfix windowing-OFDM, time-interleaved block-windowed burst OFDM, and weighted overlap and add OFDM [354]. Since CP-OFDM has established its status in mobile networks, competing technologies aim to overcome its shortcomings. However, considering individual waveforms, each of these technologies has its own pros and cons, but there is no single method that fits well to all 6G scenarios. Consequently, multiple waveforms must be utilized in ultra-heterogeneous 6G networks.
@@
IX-C2 ADVANCED MODULATION AND CODING METHODS